an experiment consists of selecting 3 light bulbs. The selected bulbs are labeled D for defective or N for non-defective.
a)list the associated sample space
b)List the event, E, the first bulb selected is defective and the last bulb
selected is non-defective.
c)What is the probability of event E?

Answer by Fombitz(32388)

There are 8 total outcomes (2^3).
a)
(N N N)
(N N D)
(N D N)
(N D D)
(D N N)
(D N D)
(D D N)
(D D D)
.
.
.
b)
E=(D N N) (D D N) 2 outcomes.
.
.
.
c)P(E)=2/8=1/4